OpenStudy (anonymous):

\[y''-4y'+y=0\] auxiliary equation \[r^2-4r+1=0\] woow I can't factor it?

OpenStudy (unklerhaukus):

try the quadratic formula

OpenStudy (anonymous):

ah ok...sorry kinda rusty on algebra...hihi

OpenStudy (unklerhaukus):

\[r=\frac{-(-4)\pm\sqrt{(-4)^2-4\times(1)\times(1)}}{2\times(1)}\]

OpenStudy (anonymous):

\[r=\frac{4\pm\sqrt{12}}{2}\]

OpenStudy (unklerhaukus):

can you simplify that square root a little bit?

OpenStudy (anonymous):

yes sir! \[r={2\pm\sqrt3}\]

OpenStudy (anonymous):

brb

OpenStudy (unklerhaukus):

nice, now should we check this somehow \[\left(r-(2+\sqrt 3)\right)\left(r-(2-\sqrt 3)\right)=0\]\[r^2-r(2-\sqrt 3)-r(2+\sqrt 3)+(2+\sqrt 3)(2-\sqrt 3)=0\]\[r^2-4r+(4+2\sqrt 3-2\sqrt 3-3)=0\]\[r^2-4r+1=0\] which is what we started with, good

OpenStudy (anonymous):

nice! ok so we then have>>>>> \[y=c_1e^{(2+\sqrt3)x}+c_2e^{(2-\sqrt3)x}\]

OpenStudy (unklerhaukus):

\[\Large\color{red} \checkmark\]Great work!

OpenStudy (unklerhaukus):

can you also write that with sines and cosines?

OpenStudy (anonymous):

oh my

OpenStudy (anonymous):

show me. In what way would I do that with sines and cosines?

OpenStudy (anonymous):

Let's do that tomorrow. It's almost 1am here, and I've gotta sleep. Thanks @UnkleRhaukus !!!!

OpenStudy (unklerhaukus):

I was thinking of a slightly harder problem if you had\[r=2\pm\sqrt{-3}\]\[r=2\pm i\sqrt{3}\]\[y(x)=c_1e^{(2+i\sqrt3)x}+c_2e^{(2-i\sqrt3)x}\] \[=e^{2x}\left(c_1e^{i\sqrt3}+c_2e^{-i\sqrt 3}\right)\] now using \[e^{i\theta}=\cos\theta +i\sin\theta\] we get \[y(x)=e^{2x}\left(\cos \sqrt 3x+\sin \sqrt 3x\right)\], but your question didn't have complex solutions so leave it as you had it